OpenStudy (anonymous):

(07.02 MC) What is the value of x in the equation 4(2x + 10) = 0? (1 point) –5 –7 5 7

OpenStudy (anonymous):

@Spring98

Mehek (mehek14):

\(4(2x + 10) = 0\\ 4*2x + 4*10 = 0\)

Mehek (mehek14):

what is \(4*2x=\)

OpenStudy (anonymous):

would you divide?? 2/4

OpenStudy (anonymous):

so x=2

OpenStudy (spring98):

x=-5

Mehek (mehek14):

\(4*2x = 8x\)

Mehek (mehek14):

\(4*10 = ?\)